OpenStudy (anonymous):

What is the equation of a parabola with a focus of (8, –4) and a directrix of y = 2?

OpenStudy (anonymous):

go to mathway it mat help you

OpenStudy (anonymous):

@akshay1234 I already have reviewed all that information . I don't understand it .

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

i am unable to do

OpenStudy (anonymous):

@annas

OpenStudy (anonymous):

@a1234 @geekfromthefutur @hidy @lululife @minnie♥mouse @mathstudent55 @Mel98 @Zinda

hero (hero):

Input points (8,-4) and (x, 2) into this formula: \[\sqrt{(x - x_1)^2 + (y - y_1)^2} = \sqrt{(x - x_2)^2 + (y - y_2)^2}\]

hero (hero):

\[\sqrt{(x - 8)^2 + (y - (-4))^2} = \sqrt{(x - x)^2 + (y - 2)^2}\] Afterwards, square both sides: \((x - 8)^2 + (y + 4)^2 = (x - x)^2 + (y - 2)^2\)

OpenStudy (anonymous):

now Do I divide by the GCF ?

hero (hero):

\(x^2 - 16x + 64 + y^2 + 8y + 16 = y^2 - 4y + 4\)

hero (hero):

Then keep simplifying: \(x^2 - 16x + 64 + 8y + 16 = -4y + 4\) \((x - 8)^2 + 12 = -12y\) \(-\dfrac{(x - 8)^2}{12} + 1 = y\) Something like that

OpenStudy (anonymous):

thats not a answer choice .

hero (hero):

What are the answer choices?

OpenStudy (anonymous):

y = -1/3(x-8)^2 -1 y = -1/3 (x+8)^2 -1 y = -1/12 (x-8)^2 - 1 y = -1/12 (x+8)^2 -1

hero (hero):

That's the same thing except I made one tiny mistake

hero (hero):

\(-\dfrac{(x - 8)^2}{12} - 1 = y\)

hero (hero):

That's the same as \(-\dfrac{1}{12}(x - 8)^2 - 1 = y\)

OpenStudy (anonymous):

would it be the same answer if y was on the other side of the equation ?

hero (hero):

Yes. In general If both sides are equal you can swap the information on both sides without changing the fact that they are equal: \(x = a\) is the same as \(a = x\)

OpenStudy (anonymous):

Find the equation of the parabola with a focus (2, 3) and directrix of y = 2.

OpenStudy (anonymous):

@Hero

hero (hero):

Follow the same method I demonstrated above.

hero (hero):

Plug the points into the formula, then simplify until you have isolated y.

OpenStudy (anonymous):

(x-2) ^2 + (y-3)^2 = (x-x)^2 + (y-2)^2 then you get .. x^2 - 4x + 4 + 2y + 4 = 3y+3 iS THAT RIGHT SO FAR ?

hero (hero):

You inserted the points correctly, but you made mistakes with the simplification.

OpenStudy (anonymous):

can you help me ?

OpenStudy (anonymous):

@Hero

hero (hero):

Try to figure out what mistakes you made. If necessary start over from the beginning.

OpenStudy (anonymous):

was all the simplifying wrong ? it all looks correct to me .

OpenStudy (anonymous):

that doesn't seem right the choices are : y = x^2/2 + 2x - 9/2 y = x^2/2 + 2x + 9/2 y = x^2/2 - 2x - 9/2 y = x^2/2 - 2x + 9/2

hero (hero):

Hang on

OpenStudy (anonymous):

Ok.

hero (hero):

Double check to make sure you posted the correct information to begin with. Post it again just to be sure.

hero (hero):

Actually, I thought about it. What I put was correct, just not in the form your problem wants.

OpenStudy (anonymous):

can you help me put it in that form ?

hero (hero):

\((x-2) ^2 + (y-3)^2 = (x-x)^2 + (y-2)^2\) \(x^2 - 4x + 4 + y^2 - 6y + 9 = y^2 - 4x + 4\) \(x^2 - 4x - 2y + 9 = 0\) \(x^2 - 4x + 9 = 2y\) \(\dfrac{x^2 - 4x + 9}{2} = y\) \(\dfrac{x^2}{2} - 2x + \dfrac{9}{2} = y\)

OpenStudy (anonymous):

@Hero can you help me with one more ? I can do it myself I just need you to check my work. I can't afford to miss any problems.

hero (hero):

Okay

OpenStudy (anonymous):

Find the equation of the parabola with focus (5, 1) and directrix y = -1.

OpenStudy (anonymous):

so far I got (x-5)^ + (y-1)^2 = (x-x)^2 + (y-2)^2 x^2 - 10x + 25 + y^2 -1y + 1 = y^2 - 10x + 25

hero (hero):

The only part that is correct so far is x^2 - 10x + 25

OpenStudy (anonymous):

I'm just copying how we did the last two problems . How am I still getting these wrong ?!

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Hero This is timed .. please help me .. :(

OpenStudy (anonymous):

@akshay1234 help please ?

OpenStudy (anonymous):

what are you trying to find?

OpenStudy (anonymous):

is this the problem What is the equation of a parabola with a focus of (8, –4) and a directrix of y = 2?

OpenStudy (anonymous):

Find the equation of the parabola with focus (5, 1) and directrix y = -1.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

ok do me a favour and repost this question so i don't have to keep scrolling down it is not very hard, we can do it in two or three steps